(Bonus: +10 pts.) Given: f(0) = cos 0 +, determine: a. The inverse function, f-10) b. The graph of the inverse function

The subject is not physic I got it by mistake,,,
it is Trigonometry
Answer #1

mil fca) = Couso + let f Coaso + = caso + I Now Swap the variables a= cos(t) + II coch (f) = a- f = cos! (0-1) Now replace of